MPU Meeting Tuesday 13th January 2009

Buildtools Project

Unsubmitted CMakeLists.txt files should perhaps be spotted and a warning given to the user - Stephen will consider it.

Power Management Project

Chris has written a man page cum design document and will distribute it for comments.

LCFG Component Testing Project

There's no time for this at the moment so we propose to stall it.

Miscellaneous Development

Operational

dhcp header
Some MPU servers were badly affected by last week's network instability because they were still using the dhcp.h header. This had been meant to be added temporarily to aid the move from KB. For now we'll remove the dhcp.h header, then reboot the affected machines as convenient.

Server reboots
we've now rebooted all the MPU servers to get the latest security and openafs fixes. We're concerned about the tardiness of vital server reboots though, and would like to find a way of getting important fixes onto key servers more quickly. One possible solution might involve using autoreboot to reboot key login servers at fixed weekly "at risk" times. Stephen will take the issue to the Operational Meeting.

Virtualisation
Two servers are now installed. The rest of the work was stalled waiting for SFPs but some have now been ordered and should arrive soon. The two most likely server virtualisation solutions will be chosen and tried out and compared.

New Inventory
Stephen needs to clear out the old sysinfo resources.

Fibre Channel kernel problems
a lack of SFPs has meant no new developments on this front, but we should have some in time to enable us to look at the problem next week.

Roaming DIY DICE on VMware
this is now documented, and Carol will it out.

Dell Optiplex 760
the first thing to report is that Dell's cheap read-only CD drive is really crap and doesn't read CDs very well. We'll try 5.3 Β when we get a copy of it.

Dual path/controller with new storage arrays
Alastair and Craig managed some of the trials: all was well as far as it went. They didn't manage to do the dual path, dual controller trials (temporary shortage of SFPs) and unfortunately there's now no possibility of doing them with the new storage arrays in the Forum as they now have to go into service as soon as possible. We'll try them with new storage arrays at KB when we have them. We now need to identify which servers to move to dual path, then move them.

Installbase/lcfg package lists
Stephen has tidied these up. The xinetd component has moved to an options header. Doing this move was complicated greatly by some servers having most of their configuration in live headers rather than release-controlled ones, making it hard to synchronise changes. We encourage all server admins to put as much server configuration as they can under release control.

Server decant from AT
it was suggested that Chris talk to Craig and maybe volunteer to move mousa. (After the meeting Chris did this and Craig liked the idea.)

bugs.lcfg.org
Chris will put bugzilla onto dresden this week. Once it's up it'll be empty of bugs, and we'll then decide what criteria to use when choosing which bugs to move from Informatics bugzilla. It should be possible to move them across in an automated bulk manoeuvre. Stephen will look at and compare his and Tim's iFriend Apache support code to check for clashes and to work out the best way to add iFriend support, which may not work as is with bugs.lcfg.org.
bradford and arbirlot
Chris has decommissioned both. He had a moan about the inventory having the wrong location data for arbirlot but apparently this is changeable with invedit not just automatically derived from switch info. Chris will get invedit access.

Default partition sizes
only Chris replied to Stephen's consultation. There's an issue over how to deal with older kit with smaller disks but it's easily solved, and Tim agrees with Stephen's thoughts on a solution. Stephen will consult COs again.

webdav and websvn
webdav for svn is done and websvn is in progress.

nvidia header changes
mostly done. Stephen now needs to get the headers in place. There's a new 180 series coming out too.

This Week

Alastair will:

  • Order Fibre Channel Cards
  • Work on Virtualisation stuff
  • Get REL 5.3 Β ISOs
  • Arrange a talk date
  • Think about Personal development topics

Chris will:

  • Power management project
  • Think about Personal development topics
  • bugs.lcfg.org
  • Move _mousa_
  • Look at packaging guidelines
  • Get invedit access
  • Remove dhcp.h from servers

Stephen will:

  • Think about Personal development topics
  • svn/webdav and websvn
  • sysinfo clearout
  • Complete nvidia changes

-- ChrisCooke - 13 Jan 2009

Topic revision: r4 - 20 Jan 2009 - 15:39:53 - StephenQuinney
 
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback
This Wiki uses Cookies